INFRARED SPECTRA OF (CO_{2})_{2}-OCS COMPLEX: INFRARED OBSERVATION OF TWO DISTINCT BARREL-SHAPED ISOMERS

J. NOROOZ OLIAEE, M. DEHGHANY, F. MIVEHVAR, \underline{N. MOAZZEN-AHMADI}, Department of Physics and Astronomy, University of Calgary, Calgary, AB T2N 1N4, Canada; A.R.W. MCKEL-LAR, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, ON K1A 0R6, Canada.

Spectra of (CO_{2})_{2}-OCS complex in the region of the OCS ν_{1} fundamental ( ∼2062 cm^{−1}) are observed using a tunable diode laser to probe a pulsed supersonic slit jet expansion. A previous microwave study of the complex by Peebles and Kuczkowskia ^{a} gave a distorted triangular cylinder. The geometerical disposition of the three dimer faces of this trimer are quite similar to the slipped CO_{2} dimer, the lowest energy form of OCS-CO_{2} (isomer a), also observed and analyzed in the microwave region, ^{b} and the higher energy form of OCS-CO_{2} (isomer b), first observed by our group in the infrared region. ^{c} Here we report the observation and analysis of two infrared bands, corresponding to two distinct isomers of the (CO_{2})_{2}- OCS complex. A band around 2058.8 cm^{−1} was assigned to isomer I, which is the same as that studied previously by microwave spectroscopy. A second band around 2051.7 cm^{−1} was assigned to a higher energy isomer of the complex, isomer II, has not been observed previously, but expected on the basis of ab initio calculations. ^{d} Approximate structural parameters for this new isomer were obtained by means of isotopic substitution. In contrast to isomer I, the geometerical disposition of the faces containing OCS and CO_{2} in isomer II are similar to isomer b of the OCS-CO_{2} complex.

aS.A. Peebles and R.L. Kuczkowski, J. Chem. Phys. 109, 5277 (1998).

bS. E. Novick, R. D. Suenram, and F. J. Lovas, J. Chem. Phys. 88, 687 (1988).

cM. Dehghany, J. Nooroz Oliaee, M. Afshari, N. Moazzen-Ahmadi, and A.R.W. McKellar, J. Chem. Phys. 130, 224310 (2009).

dH. Valdés and J. A. Sordo, Int. J. Comput. Chem. 23, 444 (2002).